If the product of a number and -7 is subtracted from the number, the result is 6 more than the number. Find the number?
Answer by Boreal(15235)

Let x= the number
x-(-7x)=x+6
x+7x=x+6
7x=6
x=(6/7) ANSWER
the product of (6/7)and (-7), which is -6, is subtracted from (6/7)
(6/7)+6=6 6/7, which is 6 more.